# Monodromy Defects in Maximally Supersymmetric Yang-Mills Theories from Holography

### v1

@@ Abstract
We study three Type II supergravity solutions which are holographically dual to codimension-2 supersymmetric defects in $(p+1)$-dimensional SU(NN) maximally supersymmetric Yang-Mills theories ($p=2,3,4$). In all of these cases, the defects have a non-trivial monodromy for the maximal abelian subgroup for the SO(9p9-p) R-symmetry. Such solutions are obtained by considering branes wrapping spindle configurations, changing the parameters (which alters the coordinate domain), and imposing suitable boundary conditions. We provide a prescription to compute the entanglement entropy of the effective theory on the defect. We find the resulting quantity to be proportional to the free energy of the ambient theory. A similar analysis is performed for the D5-brane wrapping a spindle, but we find that changing the coordinate domain does not lead to a defect solution, but rather to a circle compactification.
@@ 1 Introduction
A key element of our analysis is that in brane frame [32, 33], the near horizon metric of a Dpp-brane (for $p\neq 5$) is AdS. Although there is no conformal symmetry, as the Dilaton depends on the AdS radial coordinate, the fact that the metric in this frame is AdS, allows us to extend the same techniques used in the conformal cases mentioned above to the ones studied here. The case $p=5$ is analysed separately.
@@ 1.1 Brief Comments on Monodromy Defects
$\mathcal{A}_{R}=\mu_{R}\,\mathrm{d}z,\quad\mathcal{A}_{F_{\hat{I}}}=\mu_{\hat{I}}\,\mathrm{d}z,$ (1.2)
@@
$D_{\mu}\Psi\rightarrow\mathcal{D}_{\mu}\Psi=D_{\mu}\Psi-i\mathcal{R}[\Psi](\mathcal{A}_{R})_{\mu}\Psi-i\mathcal{F}_{\hat{I}}[\Psi](\mathcal{A}_{F_{\hat{I}}})_{\mu}\Psi,$ (1.3)
@@
Since these gauge fields are singular at $\hat{\rho}=0$, the SO($1,p$) Lorentz symmetry is broken [^2] If both the ambient theory and the defect also preserve conformal symmetry, as is the case for $p=3$, then it is broken as $\rightarrow SO(2,p12,p-1)$ $\times$ SO(2). to SO(1,p21,p-2) $\times$ SO(2), as expected for a codimension-2 defect. The effect of this background gauge field is to make fields charged under these symmetries pick-up a non-trivial monodromy around the defect. Consider for example the gaugino ($\lambda$), which is only charged under the U(1)R. Under $z\rightarrow z+2\pi$, it will pick up an extra factor of $e^{-i\mathcal{R}[\lambda]}$.
@@
We then perform a Weyl transformation $\mathrm{d}s^{2}_{p+1}=\hat{\rho}^{2}\mathrm{d}\tilde{s}^{2}_{p+1}$. In the $p=3$ case, the SYM lagrangian does not change [^3] By this, we mean the SYM lagriangian in a generic curved space, which has the spacetime covariant derivative for the fermions and the $R\Phi^{2}$ coupling. This is the case since the Weyl rescaling (1.4) changes the background where the theory is formulated from a flat to a curved one., since the theory is conformally invariant. In the non-conformal cases, the Weyl transformation does not leave the SYM action invariant. The relevant term we are interested in is the Yang-Mills kinetic term, which after the Weyl rescaling reads
@@
so that the would be gauge coupling depends on the AdSp coordinate $\hat{\rho}$. We show below that this effective gauge coupling is matched by holography. We note that, in the conformal case, $\text{AdS}_{p}\times\text{S}^{1}$ precisely realises the spacetime symmetries of the defect, while in the non-conformal case, the factor of $\hat{\rho}^{p-3}$ break the $\text{AdS}_{p}$ isometries to SO(1,p21,p-2), matching the discussion above.
@@ 2 Dpp-branes and the Dual Frame
In this section we review the Type II backgrounds holographically dual to (pp+1)-dimensional maximally supersymmetric SU($N$) Yang-Mills theories. These correspond to the decoupling limit of a stack of NN Dpp-branes. It is convenient to rewrite these solutions using the change of coordinates of [32, 33] that makes the metric conformally $\text{AdS}_{p+2}\times\text{S}^{8-p}$. It is this fact that allows us to generalize the study of codimension-2 monodromy defects in conformal theories to the maximally supersymmetric Yang-Mills theories in dimensions in which the theory is not conformally invariant. We also discuss a change of coordinates that matches the asymptotic behaviour of the monodromy defect solutions studied in the next section.
@@
$\lambda_{\text{eff}}=g^{2}_{\text{YM}}\,NE^{p-3},\qquad g^{2}_{YM}=g_{s}(2\pi)^{p-2}(\alpha^{\prime})^{\frac{p-3}{2}}\\$ (2.7)
@@
To obtain the background dual to a codimension-2 monodromy defect in maximally supersymmetric Yang-Mills in $(p+1)$ dimensions, we can extend the procedure of [30, 22, 23, 26, 24, 25] to non-conformal branes. This is, we start by considering the solutions from [34] of Dpp-branes ($p=2,4,5$) wrapped on spindles [^10] The $p=3$ case was studied in [24], but we briefly review it for consistency.. Then, by considering a different range of the spindle coordinates, we can interpret the solution as a codimension-2 monodromy defect. We discuss this in detail in Section 3 for $p=2,3,4$ and Section 4 for $p=5$.
@@ 3.1 General Aspects of the Solutions and Boundary Conditions
We start by considering the solutions of Dpp-branes wrapped on a spindle of [34]. These are solutions of a U(1)r truncation of the $(p+2)$-dimensional maximal SO(9p9-p) gauged supergravity, here $\mathfrak{r}$ is the rank of SO(9p9-p), obtained by reducing the 10D theory [^11] For $p=2$, that is, the reduction on $\text{S}^{6}$, the resulting 4D maximally supersymmetric theory has ISO(7) as gauge group. Here we consider only the compact SO(7) subgroup. on $\text{S}^{8-p}$. The bosonic field content of these lower dimensional U(1)r theories are just the metric, $\mathfrak{r}$ U(1) gauge fields and a proper number of scalars depending on $p$. The metric of the solutions under consideration takes the form [^12] We perform a change of coordinates of the form $y\rightarrow y^{c}$ of the solutions of [34] and a redefinition of $H$ and $P$ such that all the backgrounds behave in the way presented here.
@@
where $I=1,...,\mathfrak{r}$ and $\Phi^{(p)}$ denotes the sets of scalars of the ($p+2$)-dimensional theory. The explicit form of $F$, $H$, $P$, $a^{I}$ and the scalars depends explicitly on $p$. We note that the supersymmetric solutions of the U(1)r theories under consideration are characterized by $2\mathfrak{r}$ integration constants $(q_{I},\alpha_{I})$.
@@
  • We allow the metric and scalars to be singular at $\rho=0$ and $\rho\rightarrow+\infty$, since those are present in the vacuum solutions (2.14a)-(2.14c).
@@
The first configuration we study corresponds to the D2-brane wrapping spindle of [34], which is a solution of the 4D U(1)3 truncation of the maximal ISO(7) gauged supergravity [40]. Here, we slightly change conventions to match the form of the solutions in (3.1), such that the vacuum solution (in brane frame) takes the form (2.14a)-(2.14b). The solution takes the form
$\begin{split}\mathrm{d}s^{2}&=\frac{2^{14/3}}{3^{7/3}g^{7/3}}y^{1/3}H^{1/2}\rho^{1/3}\left(-\rho^{2}\mathrm{d}t^{2}+\frac{\mathrm{d}\rho^{2}}{\rho^{2}}+\frac{\mathrm{d}y^{2}}{P}+\frac{P}{H}n^{2}\mathrm{d}z^{2}\right),\\[5.69054pt] A^{I}&=\left(\alpha_{I}+\kappa\frac{4}{3g}\sqrt{1-\frac{\nu}{q_{I}}}\frac{q_{I}}{h_{I}}n\right)\mathrm{d}z,\\[5.69054pt] e^{\phi_{0}}&=\frac{2^{4/3}}{3^{2/3}g^{2/3}}y^{2/3}\rho^{2/3},\qquad e^{\phi_{I}}=\frac{2^{2/3}}{3^{1/3}g^{1/3}}\frac{\rho^{1/3}\sqrt{H}}{y^{1/3}h_{I}},\end{split}$ (3.9)
@@
We have also included the BPS-parameter $\nu$ that completely breaks supersymmetry. Besides $\nu$, the configuration is characterized by six integration constants $(\alpha_{I},q_{I})$. Also, $n$ is a conical singularity parameter (recall $z\sim z+2\pi$), while $\kappa=\pm 1$. The vacuum solution is obtained by setting $(\alpha_{I},q_{I})=0$ and $n=1$. We provide the uplift of to Type IIA in Appendix A.5.
In what follows we focus on the supersymmetric defect solutions, so we set $\nu=0$. We also fix $\kappa=1$ for convenience, as it can always be set to this value by $z\rightarrow-z$.
@@
$q_{I}=-\frac{3g\mu^{I}y_{\text{core}}^{4/3}}{3g\mu^{I}+4n},\qquad y_{\text{core}}=\frac{\sqrt{3g\mu^{1}+4n}\sqrt{3g\mu^{2}+4n}\sqrt{3g\mu^{3}+4n}}{8n^{3/2}}.$ (3.11)
Then, the next step is to impose a smoothly shrinking geometry at $y=y_{\text{core}}$ (3.4). This, combined with (3.11) leads to a constraint for the background gauge fields $\mu_{I}$
$g\mu^{1}+g\mu^{2}+g\mu^{3}=2(n-1).$ (3.12)
Using (A.29), we can rewrite this constraint and (3.11) in terms of the background gauge field for the flavour symmetries $(\mu_{F_{1}},\mu_{F_{2}})$ and the R-symmetry $\mu_{R}$. First, we note that (3.12) only fixes the R-symmetry background gauge field
$\mu_{R}=\frac{2}{g}(n-1).$ (3.13)
@@
$y_{\text{core}}=\frac{\sqrt{2(n+1)-g\mu_{F_{1}}-2g\mu_{F_{2}}}\sqrt{2(n+1)-g\mu_{F_{1}}+g\mu_{F_{2}}}\sqrt{2(n+1)+2g\mu_{F_{1}}+g\mu_{F_{2}}}}{8n^{3/2}}.$ (3.14)
@@
$y_{\text{core}}=\frac{(1+n)^{3/2}}{2^{3/2}n^{3/2}},\qquad q_{1}=\frac{n^{2}-1}{4n^{2}},\qquad\alpha_{1}=\mu^{1}=\frac{2}{3g}(1-n).$ (3.15)
@@
Setting some of the gauge fields to either be zero leads to more supersymmetry being preserved (at the 10D level). This can be understood simply by looking at the uplift in Appendix (A.5), as a trivial gauge fields restores a subgroup of the isometries of the internal manifold.
We focus on the simplest case $A^{2}=A^{3}=0$. This restores an $SO(4)=SU(2)\times SU(2)$ worth of isometries in the internal manifold, so we have four times as much supersymmetry. In this case, the parameters defining the solution read
$y_{\text{core}}=\frac{\sqrt{3-n}}{\sqrt{2}\sqrt{n}},\qquad q_{1}=\frac{3(n-1)}{2^{2/3}(3-n)^{1/3}n^{2/3}},\qquad\alpha_{1}=\mu^{1}=\frac{2}{g}(1-n).$ (3.16)
@@
In this section we briefly review the 5D supergravity background dual to a codimension-2 defect on 4D $\mathcal{N}=4$ SU(NN) SYM. This solution was obtained in [41] and recently studied in [42, 43, 44] in the context of black hole near horizon geometries and spindle compactifications. The extension of this solution that allows to have a codimension-2 superconformal defect interpretation was investigated in [24], where this solution was deeply analysed. We limit ourselves to presenting the solution and reviewing the main features, since it completes our discussion for codimension-2 defects in Dpp-brane theories.
The background is a solution of the 5D U(1)3 gauged supergravity. The bosonic fields are the metric, three gauge fields $A^{I}$ ($I=1,2,3$) and two scalars $\phi_{1}$ and $\phi_{2}$. See Appendix A.4 for details. In the conventions of (3.1), supergravity background reads
@@
Here, the $\text{AdS}_{3}$ factor is of unit radius metric and SO(2,2) is an exact isometry of the configuration. The functions $h_{I}$, $H$ and $P$ depend only on $y$ and are given by
@@
As before, $\nu$ is a non-supersymmetric parameter and $\kappa=\pm 1$. Furthermore, using the uplift in Appendix A.4 and imposing flux quantization we find $g=1/\ell$, with $\ell$ given by $p=3$ in (2.2).
We now study defect boundary conditions for the supersymmetric solution, so we set $\nu=0$. In this case, the configuration preserved $\mathcal{N}=(0,2)$ supersymmetry [20]. Also, for convenience we take $\kappa=1$. Using (3.7) and (3.8) we find
@@
$g\mu_{1}+g\mu_{2}+g\mu_{3}=(n-1),\quad\Rightarrow\quad\mu_{R}=\frac{1}{g}(n-1),$ (3.20)
@@
where we included the non-BPS parameter $\nu$. Here, we have made a change of variables with respect to [34], such that setting $q_{I}=\alpha_{I}=\nu=0$, using the uplift in Appendix A.3 and moving to the dual frame, we recover the vacuum written in the form of (2.14a). Before proceeding to the study of defect boundary conditions, let us comment that this solution can be lifted to the 7D U(1)2 AdS gauged supergravity that has been deeply studied as spindle-compactification [45, 44] and defect [24, 22, 23]. Focusing on the defect interpretation, it was shown in [24, 22, 23] that it is possible define values of the integration constants such that the solution can be interpreted as a codimension-2 monodromy defect on a 6D dual CFT. Since this proves that it is possible to obtain a defect solution in the 6D supergravity theory, we briefly repeat the analysis in our parametrization for completeness. Below, we choose $\kappa=1$ for convenience, and we focus on the supersymmetric solution, so we set $\nu=0$.
@@
In the case we set to zero one of the gauge fields, say $A^{2}=0$, the 10D the number of supercharges preserved is doubled. This was already observed at the level of the 7D solution [23] (recall that the 6D solution under consideration can be obtained by dimensional reduction from the 7D one). The main difference being that in 6D there are no AdS isometries, so there are not superconformal killing spinors. In this case, we have

### v2

@@ Abstract
We study three Type II supergravity solutions holographically dual to codimension-2 supersymmetric defects in $(p+1)$-dimensional SU(NN) maximally supersymmetric Yang-Mills theory ($p=2,3,4$). In all of these cases, the defects have a non-trivial monodromy for the maximal abelian subgroup for the SO(9p9-p) R-symmetry. Such solutions are obtained by considering branes wrapping spindle configurations, changing the parameters (which alters the coordinate domain), and imposing suitable boundary conditions. We provide a prescription to compute the entanglement entropy of the effective theory on the defect. We find the resulting quantity to be proportional to the free energy of the ambient theory. A similar analysis is performed for the D5-brane wrapping a spindle, but we find that changing the coordinate domain does not lead to a defect solution, but rather to a circle compactification.
@@ 1 Introduction
A key element of our analysis is that the brane frame metric of the decoupling limit of a stack of Dpp-brane (for $p\neq 5$) is AdS [70, 57]. Although there is no conformal symmetry, as the Dilaton depends on the AdS radial coordinate, the fact that the metric in this frame is AdS, allows us to extend the same techniques used in the conformal cases mentioned above to the ones studied here. The case $p=5$ is analysed separately.
@@ 1.1 Brief Comments on Monodromy Defects
$\mathcal{A}_{\text{R}}=\mu_{\text{R}}\,\mathrm{d}z,\quad\mathcal{A}_{F_{\hat{I}}}=\mu_{\hat{I}}\,\mathrm{d}z,$ (1.2)
@@
$D_{\mu}\Psi\rightarrow\mathcal{D}_{\mu}\Psi=D_{\mu}\Psi-i\mathcal{R}[\Psi](\mathcal{A}_{\text{R}})_{\mu}\Psi-i\mathcal{F}_{\hat{I}}[\Psi](\mathcal{A}_{F_{\hat{I}}})_{\mu}\Psi,$ (1.3)
@@
[…]
@@
We then perform a Weyl transformation $\mathrm{d}s^{2}_{p+1}=\hat{\rho}^{2}\mathrm{d}\tilde{s}^{2}_{p+1}$. In the $p=3$ case, the SYM lagrangian does not change [^3] By this, we mean the SYM lagriangian in a generic curved space, which has the spacetime covariant derivative for the fermions and the $R\Phi^{2}$ coupling. This is the case since the Weyl rescaling (1.4) changes the background where the theory is formulated from a flat to a curved one., since the theory is conformally invariant. In the non-conformal cases, the Weyl transformation does not leave the SYM action invariant. However, due to the generalized conformal structure preserved by maximally supersymmetric Yang-Mills [57], the Weyl rescaling leads to a position dependent gauge coupling. For example, focusing on the Yang-Mills kinetic term, which after the Weyl rescaling reads
@@
the would be gauge coupling depends on the AdSp radial coordinate $\hat{\rho}$. We show below that this effective gauge coupling is matched by holography.
In the conformal case, $\text{AdS}_{p}\times\text{S}^{1}$ precisely realises the spacetime symmetries of a codimension-2 conformal defect, while in the non-conformal case, the factor of $\hat{\rho}^{p-3}$ break the $\text{AdS}_{p}$ isometries to SO(1,p21,p-2), matching the discussion above.
[…]
@@ 2 Dpp-branes and the Dual Frame
In this section we review the Type II backgrounds holographically dual to (pp+1)-dimensional maximally supersymmetric SU($N$) Yang-Mills theory. These correspond to the decoupling limit of a stack of NN Dpp-branes. It is convenient to rewrite these solutions using the change of coordinates of [70, 57] that makes the metric conformally $\text{AdS}_{p+2}\times\text{S}^{8-p}$. It is this fact that allows us to generalize the study of codimension-2 monodromy defects in conformal theories to the maximally supersymmetric Yang-Mills theories in dimensions in which the theory is not conformally invariant. We also discuss a change of coordinates that matches the asymptotic behaviour of the monodromy defect solutions studied in the next section.
@@
$\lambda_{\text{eff}}=g^{2}_{\text{YM}}\,NE^{p-3},\qquad g^{2}_{\text{YM}}=g_{s}(2\pi)^{p-2}(\alpha^{\prime})^{\frac{p-3}{2}}\\$ (2.7)
@@
To obtain the background dual to a codimension-2 monodromy defect in maximally supersymmetric Yang-Mills in $(p+1)$ dimensions, we can extend the procedure of [20, 49, 48, 16, 28, 29] to non-conformal branes. This is, we start by considering the solutions from [15] of Dpp-branes ($p=2,4,5$) wrapped on spindles [^10] The $p=3$ case was studied in [20, 49, 48, 16, 28, 29], but we briefly review it for consistency.. Then, by considering a different range of the spindle coordinates, we can interpret the solution as a codimension-2 monodromy defect. We discuss this in detail in Section 3 for $p=2,3,4$ and Section 4 for $p=5$.
@@ 3.1 General Aspects of the Solutions and Boundary Conditions
We start by considering the solutions of Dpp-branes wrapped on a spindle of [15]. These are solutions of a U(1)r truncation of the $(p+2)$-dimensional SO(9p9-p) gauged supergravity, here $\mathfrak{r}$ is the rank of SO(9p9-p), obtained by reducing the 10D theory [^11] For $p=2$, that is, the reduction on $\text{S}^{6}$, the resulting 4D gauged supergravity has ISO(7) as gauge group. Here we consider only the compact SO(7) subgroup. on $\text{S}^{8-p}$. The bosonic field content of these lower dimensional U(1)r theories are: the metric, $\mathfrak{r}$ U(1) gauge fields and a proper number of scalars whose number depends on $p$. The metric of the solutions under consideration takes the form [^12] We perform a change of coordinates of the form $y\rightarrow y^{c}$ of the solutions of [15] and a redefinition of $H$ and $P$ such that all the backgrounds behave in the way presented here.
@@
where $I=1,...,\mathfrak{r}$ and $\Phi^{(p)}$ denotes the set of scalars of the ($p+2$)-dimensional theory. The explicit form of $F$, $H$, $P$, $a^{I}$ and the scalars depends explicitly on $p$ (we explicitly show the solutions below). We note that the supersymmetric solutions of the U(1)r theories under consideration are characterized by $2\mathfrak{r}$ integration constants $(q_{I},\alpha_{I})$.
@@
  • We allow the metric and scalars to be singular at $\rho=0$ and $\rho\rightarrow+\infty$, since these singularities are already present in the vacuum solutions (2.14a)-(2.14c).
@@
The first configuration we study corresponds to the D2-brane wrapping a spindle of [^13] We use conventions such that the gauge coupling $g_{\text{here}}=\frac{1}{2}g_{\text{there}}$. [15], which is a solution of the 4D U(1)3 truncation of the ISO(7) gauged supergravity [53]. Here, we slightly change conventions to match the form of the solutions in (3.1), such that the vacuum solution (in brane frame) takes the form (2.14a)-(2.14b). The solution takes the form
$\begin{split}\mathrm{d}s^{2}&=\left(\frac{2}{3g}\right)^{7/3}y^{1/3}H^{1/2}\rho^{1/3}\left(-\rho^{2}\mathrm{d}t^{2}+\frac{\mathrm{d}\rho^{2}}{\rho^{2}}+\frac{\mathrm{d}y^{2}}{P}+\frac{P}{H}n^{2}\mathrm{d}z^{2}\right),\\[5.69054pt] A^{I}&=\left(\alpha_{I}+\kappa\frac{2}{3g}\sqrt{1-\frac{\nu}{q_{I}}}\frac{q_{I}}{h_{I}}n\right)\mathrm{d}z,\\[5.69054pt] e^{\phi_{0}}&=\left(\frac{2}{3g}\right)^{2/3}y^{2/3}\rho^{2/3},\qquad e^{\phi_{I}}=\left(\frac{2}{3g}\right)^{1/3}\frac{\rho^{1/3}\sqrt{H}}{y^{1/3}h_{I}},\end{split}$ (3.9)
@@
We have also included the BPS-parameter $\nu$ that completely breaks supersymmetry. Besides $\nu$, the configuration is characterized by six integration constants $(\alpha_{I},q_{I})$. Also, $n$ is a conical singularity parameter (recall $z\sim z+2\pi$), while $\kappa=\pm 1$. The vacuum solution is obtained by setting $(\alpha_{I},q_{I})=0$ and $n=1$. We provide the uplift to Type IIA in Appendix A.5.
In what follows we focus on the supersymmetric defect solutions, so we set $\nu=0$. For generic $(\alpha_{I},q_{I})$ the solution preserves two supercharges. We also fix $\kappa=1$ for convenience, as it can always be set to this value by $z\rightarrow-z$.
@@
$q_{I}=-\frac{3g\mu^{I}y_{\text{core}}^{4/3}}{3g\mu^{I}+2n},\qquad y_{\text{core}}=\frac{\sqrt{3g\mu^{1}+2n}\sqrt{3g\mu^{2}+2n}\sqrt{3g\mu^{3}+2n}}{\sqrt{8}n^{3/2}}.$ (3.11)
Then, the next step is to impose a smoothly shrinking geometry at $y=y_{\text{core}}$, as in (3.4). This, combined with (3.11) leads to a constraint for the background gauge fields $\mu_{I}$
$g\mu^{1}+g\mu^{2}+g\mu^{3}=n-1.$ (3.12)
Using (A.29), we can rewrite this constraint and (3.11) in terms of the background gauge field for the flavour symmetries $(\mu_{F_{1}},\mu_{F_{2}})$ and the R-symmetry $\mu_{\text{R}}$. First, we note that (3.12) only fixes the R-symmetry background gauge field
$g\mu_{\text{R}}=n-1.$ (3.13)
@@
$y_{\text{core}}=\frac{\sqrt{n+1-g\mu_{F_{1}}-2g\mu_{F_{2}}}\sqrt{n+1-g\mu_{F_{1}}+g\mu_{F_{2}}}\sqrt{n+1+2g\mu_{F_{1}}+g\mu_{F_{2}}}}{\sqrt{8}n^{3/2}}.$ (3.14)
@@
$y_{\text{core}}=\frac{(1+n)^{3/2}}{2^{3/2}n^{3/2}},\qquad q_{1}=\frac{n^{2}-1}{4n^{2}},\qquad\alpha_{1}=\mu^{1}=\frac{1}{3g}(1-n).$ (3.15)
@@
Setting some of the gauge fields to be zero leads to more supersymmetry being preserved (at the 10D level). This can be understood simply by looking at the uplift in Appendix (A.5), as a trivial gauge field restores a subgroup of the isometries of the internal manifold.
We focus on the simplest case $A^{2}=A^{3}=0$. This restores an SO(4) $\cong$ $\times SU(2)$ worth of isometries in the internal manifold, so we have four times as much supersymmetry. In this case, the parameters defining the solution read
$y_{\text{core}}=\frac{\sqrt{3-n}}{\sqrt{2}\sqrt{n}},\qquad q_{1}=\frac{3(n-1)}{2^{2/3}(3-n)^{1/3}n^{2/3}},\qquad\alpha_{1}=\mu^{1}=\frac{1}{g}(1-n).$ (3.16)
@@
In this section we briefly review the 5D supergravity background dual to a codimension-2 defect on 4D $\mathcal{N}=4$ SU(NN) SYM. This solution was obtained in [60] and recently studied in [14, 52, 40] in the context of black hole near horizon geometries and spindle compactifications. The extension of this solution that allows to have a codimension-2 superconformal defect interpretation was investigated in [28], where this solution was analysed. We limit ourselves to presenting the solution and reviewing the main features, since it completes our discussion for codimension-2 defects in Dpp-brane theories.
The background is a solution of the 5D U(1)3 gauged supergravity. The bosonic fields are the metric, three gauge fields $A^{I}$ ($I=1,2,3$) and two scalars $\phi_{1}$ and $\phi_{2}$. See Appendix A.4 for details. In the conventions of (3.1), the supergravity background reads
@@
Here, the $\text{AdS}_{3}$ factor has unit radius and SO(2,2) is an exact isometry of the configuration. The functions $h_{I}$, $H$ and $P$ depend only on $y$ and are given by
@@
As before, $\nu$ is a non-supersymmetric parameter and $\kappa=\pm 1$. Furthermore, using the uplift in Appendix A.4 and imposing flux quantization we find $g=1/\ell$, with $\ell$ given by (2.2) with $p=3$.
We now study defect boundary conditions for the supersymmetric solution, so we set $\nu=0$. In this case, the configuration preserves $\mathcal{N}=(0,2)$ supersymmetry [3]. Also, for convenience we take $\kappa=1$. Using (3.7) and (3.8) we find
@@
$g\mu_{1}+g\mu_{2}+g\mu_{3}=n-1,\quad\Rightarrow\quad g\mu_{\text{R}}=n-1,$ (3.20)
@@
where we included the non-BPS parameter $\nu$. Here, we have made a change of variables with respect to [15], such that setting $q_{I}=\alpha_{I}=\nu=0$, using the uplift in Appendix A.3 and moving to the dual frame, we recover the vacuum written in the form of (2.14a). Before proceeding to the study of defect boundary conditions, let us comment that this solution can be lifted to the 7D U(1)2 AdS gauged supergravity that has been studied as spindle-compactification [39, 40] and defect [28, 49, 48]. Focusing on the defect interpretation, it was shown in [28, 49, 48] that it is possible to define values of the integration constants such that the solution can be interpreted as a codimension-2 monodromy defect on a 6D dual CFT. Since this proves that it is possible to obtain a defect solution in the 6D supergravity theory, we briefly repeat the analysis in our parametrization for completeness. Below, we choose $\kappa=1$ for convenience, and we focus on the supersymmetric solution, so we set $\nu=0$.
@@
In the case we set to zero one of the gauge fields, say $A^{2}=0$, 10D the number of supercharges preserved is doubled. This was already observed at the level of the 7D solution [48] (recall that the 6D solution under consideration can be obtained by dimensional reduction from the 7D one). The main difference being that in 6D there are no AdS isometries, so there are not superconformal killing spinors. In this case, we have